OpenStudy (anonymous):

x(x+22)=11(4x-11) would it would be (x+11)(x-2)=0 ?

OpenStudy (amistre64):

x^2+22x = 44x -121 x^2 -22x +122 = 0 (x-11)^2 = 0

OpenStudy (anonymous):

therefore, its 11?

OpenStudy (amistre64):

sounds good to me :) 11(11+22) =? 11(44-11) 11(33) =? 11(33) yep

OpenStudy (anonymous):

sweet :D thanks :D i swear i need a math tutor haha. College math sucks =/

OpenStudy (amistre64):

im taking calc3, diffy qs, and linear alg next term

OpenStudy (amistre64):

im not to matrix literate so thatll be a little tough ;)

OpenStudy (anonymous):

=O this semster its Math 95. next semester [in january] its math 137 :( looks like ima be on here a lot =/